what will be the unit digit of 77895463*85931457858

Grade:10

4 Answers

Hasan Naqvi
8 years ago
4
 
since the unit digit of product of unit digits of given no = 4 (3 * 8 = 24)
 
so the unit digit of entire multiplication will be 4.
grenade
8 years ago
4
bcoz 
3*8 lead to 4 @ unit place
Nicho priyatham
8 years ago
4
the units digit in multiplication is always the unitdigit of product of unit digits of given number
3*8=24
unit digit of 24 is 4 
Prajwal Kavad
8 years ago
4 coz 3